Please help me
Find two numbers whose sum is 7 and whose product is 10

Answer by scott8148(5895)

like 2 and 5?

using "algebra" ... x+y=7 ... xy=10

x(7-x)=10 ... 0=x^2-7x+10 ... 0=(x-5)(x-2) ... x=5, and x=2 ... so y equals the other one